OpenStudy (anonymous):

A motor scooter purchased for $1,000 depreciates at an annual rate of 15%. Write an exponential function, and graph the function. Use the graph to predict when the value will fall below $100.

OpenStudy (anonymous):

A. v(t) = 1000(0.85)t; the value will fall below $100 in about 14.2 yr. B. v(t) = 1000(0.85)t; the value will fall below $100 in about 18 yr. C. v(t) = 0.85(1000)t; the value will fall below $100 in about 14.2 yr. D. v(t) = 0.85(1000)t; the value will fall below $100 in about 18 yr.

OpenStudy (anonymous):

I got A